# Exponential separations between classical and quantum learners

### v1

@@ 5.2 Power of data
In [${}^{+}$21] the authors show how having access to data radically enhances what can be efficiently evaluated. In this section we connect the ideas from their work to the formalism we introduce in this paper. Specifically, we will discuss a family of functions inspired by [${}^{+}$21] that from their description alone cannot be efficiently evaluated classically, yet access to a few examples (i.e., evaluations of the function) allows them to be efficiently evaluated classically. This highlights an important difference between complexity-theoretic separations and learning separations, since in the latter one has to deal with the learner having access to data when proving classical non-learnability.
Consider a polynomial-depth parameterized quantum circuit $U(\theta,\vec{\phi})$ – with two types of parameters $\theta\in\mathbb{R}$ parameterizing a single gate and ϕitalic-$\vec{\phi}\in\mathbb{R}^{\ell}$ parameterizing multiple other gates – that is universal in the sense that for every polynomial-depth circuit $V$ there exists parameters ϕ*superscriptitalic-$\vec{\phi}^{*}\in\mathbb{R}^{\ell}$ such that
@@
Following an argument similar to [${}^{+}$21], due to the universality of the parameterized quantum circuit no efficient randomized classical algorithm can take as input a ϕitalic-$\vec{\phi}\in\mathbb{R}^{\ell}$ and compute the function $f_{\vec{\phi}}$ on a given point $\theta\in\mathbb{R}$ up to constant error in time $\mathcal{O}\left(\mathrm{poly}(n)\right)$, unless $\mathsf{BPP}=\mathsf{BQP}$. Intuitively, one might thus think that the concept class $\{f_{\vec{\phi}}\mid\vec{\phi}\in\mathbb{R}^{\ell}\}$ exhibits a separation between classical and quantum learners. However, it turns out that the examples given to a classical learner radically enhance what it can efficiently evaluate. In particular, given a few of evaluations of $f_{\vec{\phi}}$ for some fixed but arbitrary ϕitalic-$\vec{\phi}\in\mathbb{R}^{\ell}$, a classical learner is suddenly able to efficiently evaluate the function. To see this, note that by [NFT20] one can write the functions as
@@
As discussed in Section 3, one way to deal with the fact that data can radically enhance what can be efficiently evaluated is to ensure that the data itself is efficiently generatable. However, for the concepts discussed above, the examples are such that only a quantum computer can generate them efficiently. In other words, these functions exemplify how hard to generate data can radically enhance what a classical learner can efficiently evaluate. As discussed in Section 3, another way to deal with the fact that data can radically enhance what can be efficiently evaluated is to ensure that the concepts lie outside of $\mathsf{HeurP}/\mathsf{poly}$. However, for the case discussed above, every $f_{\vec{\phi}}$ corresponds to a function in $\mathsf{HeurP}/\mathsf{poly}$, since the coefficients $\alpha,\beta$ and $\gamma$ suffice as the advice. Finally, we note that for certain circuits one could have exponentially many terms in the Fourier series [CCL23, ${}^{+}$21], in which case it is unclear how to classically learn them.
@@ 5.3 Physically-motivated PAC learning settings with fixed hypothesis classes
As in the case of Hamiltonian learning, one might think that the above concepts are good candidates to exhibit a $\mathsf{C}_{\mathcal{H}^{\mathrm{order}}}/\mathsf{Q}_{\mathcal{H}^{\mathrm{order}}}$ separation, since the hypotheses are classically intractable and quantumly efficient to evaluate (assuming $\mathsf{BPP}\neq\mathsf{BQP}$). In fact, according to the folklore, quantum learners are most likely to have advantages for data that is “quantum-generated”, which certainly seems to also be the case here. However, as already mentioned, in the setting where the learner is constrained to output hypotheses from a fixed hypothesis class the goal is only to identify the correct hypothesis, and it is therefore not enough to just have concepts that are classically intractable. We remark that the methods of [${}^{+}$22b] also apply to phase classification, but they are more aimed at the PAC learning setting where the learner can output arbitrary hypotheses (i.e., the main goal is to predict the phase of a given physical system). In particular, their methods do not directly allow one to obtain a physically-meaningful description of the order-parameter, which is the main goal in the setting where the learner is constrained to output hypotheses from a fixed hypothesis class (which is related to the popular theme of “explainability” in machine learning).
@@ Acknowledgements
The authors thank Simon C. Marshall, Srinivasan Arunachalam, and Tom O’Brien for helpful discussions. The authors are grateful to Peter Bruin for valuable comments on the hardness of the discrete cube root for the $2^{c}$-integers in Section 3.3. This work was supported by the Dutch Research Council (NWO/ OCW), as part of the Quantum Software Consortium programme (project number 024.003.037). This work was supported by the Dutch National Growth Fund (NGF), as part of the Quantum Delta NL programme.
@@ A.3 Proof of Lemma 2
The proof strategy is analogous to the arguments in Section 2 of the supplementary material of [${}^{+}$21], where they show that $\mathsf{BPP/samp}\subseteq\mathsf{P/poly}$. Let $(L,\{\mathcal{D}_{n}\}_{n\in\mathbb{N}})~{}\in~{}\mathsf{HeurBPP/samp}$. In particular, there exist a polynomial-time classical algorithms $\mathcal{A}$ with the following property: for every $n$ and $\epsilon>0$ it holds that
@@
By applying the arguments of Section 2 of the supplementary material of [${}^{+}$21] to $\mathcal{A}$ with the bitstring $0^{\lfloor 1/\epsilon\rfloor}$ fixed as input we obtain a polynomial-time classical algorithm $\mathcal{A}^{\prime}$ with the following property: for every $n$ there exists an advice string $\alpha_{n,\epsilon}\in\{0,1\}^{\mathrm{poly(n,1/\epsilon)}}$ [^1] Note that the advice string also depends on $\epsilon$, since $0^{\lfloor 1/\epsilon\rfloor}$ was fixed as input to $\mathcal{A}$. such that for every $x\in I^{n}_{\text{correct}}(\epsilon)$:
@@
Appendix B Proof of Theorem 3

### v2

@@ 5.2 Power of data
In [HBM+21] the authors show how having access to data radically enhances what can be efficiently evaluated. In this section we connect the ideas from their work to the formalism we introduce in this paper. Specifically, we will discuss a family of functions inspired by [HBM+21] that from their description alone cannot be efficiently evaluated classically, yet access to a few examples (i.e., evaluations of the function) allows them to be efficiently evaluated classically. This highlights an important difference between complexity-theoretic separations and learning separations, since in the latter one has to deal with the learner having access to data when proving classical non-learnability.
Consider a polynomial-depth parameterized quantum circuit $U(\theta,\vec{\phi})$ – with two types of parameters $\theta\in\mathbb{R}$ parameterizing a single gate and ϕitalic-$\vec{\phi}\in\mathbb{R}^{\ell}$ parameterizing multiple other gates – that is universal in the sense that for every polynomial-depth circuit $V$ there exists parameters ϕsuperscriptitalic-$\vec{\phi}^{*}\in\mathbb{R}^{\ell}$ such that
@@
Following an argument similar to [HBM+21], due to the universality of the parameterized quantum circuit no efficient randomized classical algorithm can take as input a ϕitalic-$\vec{\phi}\in\mathbb{R}^{\ell}$ and compute the function $f_{\vec{\phi}}$ on a given point $\theta\in\mathbb{R}$ up to constant error in time $\mathcal{O}\left(\mathrm{poly}(n)\right)$, unless $\mathsf{BPP}=\mathsf{BQP}$. Intuitively, one might thus think that the concept class $\{f_{\vec{\phi}}\mid\vec{\phi}\in\mathbb{R}^{\ell}\}$ exhibits a separation between classical and quantum learners. However, it turns out that the examples given to a classical learner radically enhance what it can efficiently evaluate. In particular, given a few of evaluations of $f_{\vec{\phi}}$ for some fixed but arbitrary ϕitalic-$\vec{\phi}\in\mathbb{R}^{\ell}$, a classical learner is suddenly able to efficiently evaluate the function. To see this, note that by [NFT20] one can write the functions as
@@
As discussed in Section 3, one way to deal with the fact that data can radically enhance what can be efficiently evaluated is to ensure that the data itself is efficiently generatable. However, for the concepts discussed above, the examples are such that only a quantum computer can generate them efficiently. In other words, these functions exemplify how hard to generate data can radically enhance what a classical learner can efficiently evaluate. As discussed in Section 3, another way to deal with the fact that data can radically enhance what can be efficiently evaluated is to ensure that the concepts lie outside of $\mathsf{HeurP}/\mathsf{poly}$. However, for the case discussed above, every $f_{\vec{\phi}}$ corresponds to a function in $\mathsf{HeurP}/\mathsf{poly}$, since the coefficients $\alpha,\beta$ and $\gamma$ suffice as the advice. Finally, we note that for certain circuits one could have exponentially many terms in the Fourier series [CCL23, CGFM+21], in which case it is unclear how to classically learn them.
@@ 5.3 Physically-motivated PAC learning settings with fixed hypothesis classes
As in the case of Hamiltonian learning, one might think that the above concepts are good candidates to exhibit a $\mathsf{C}_{\mathcal{H}^{\mathrm{order}}}/\mathsf{Q}_{\mathcal{H}^{\mathrm{order}}}$ separation, since the hypotheses are classically intractable and quantumly efficient to evaluate (assuming $\mathsf{BPP}\neq\mathsf{BQP}$). In fact, according to the folklore, quantum learners are most likely to have advantages for data that is “quantum-generated”, which certainly seems to also be the case here. However, as already mentioned, in the setting where the learner is constrained to output hypotheses from a fixed hypothesis class the goal is only to identify the correct hypothesis, and it is therefore not enough to just have concepts that are classically intractable. We remark that the methods of [HKT+22b] also apply to phase classification, but they are more aimed at the PAC learning setting where the learner can output arbitrary hypotheses (i.e., the main goal is to predict the phase of a given physical system). In particular, their methods do not directly allow one to obtain a physically-meaningful description of the order-parameter, which is the main goal in the setting where the learner is constrained to output hypotheses from a fixed hypothesis class (which is related to the popular theme of “explainability” in machine learning).
@@ Acknowledgements
The authors thank Simon C. Marshall, Srinivasan Arunachalam, and Tom O’Brien for helpful discussions. The authors are grateful to Peter Bruin for valuable comments on the hardness of the discrete cube root for the $2^{c}$-integers in Section 3.3. This work was supported by the Dutch Research Council (NWO/ OCW), as part of the Quantum Software Consortium programme (project number 024.003.037). This work was also supported by the Dutch National Growth Fund (NGF), as part of the Quantum Delta NL programme. This work was also supported by the European Union’s Horizon Europe program through the ERC CoG BeMAIQuantum (Grant No. 101124342). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
@@ A.3 Proof of Lemma 2
The proof strategy is analogous to the arguments in Section 2 of the supplementary material of [HBM+21], where they show that $\mathsf{BPP/samp}\subseteq\mathsf{P/poly}$. Let $(L,\{\mathcal{D}_{n}\}_{n\in\mathbb{N}})~{}\in~{}\mathsf{HeurBPP/samp}$. In particular, there exist a polynomial-time classical algorithms $\mathcal{A}$ with the following property: for every $n$ and $\epsilon>0$ it holds that
@@
By applying the arguments of Section 2 of the supplementary material of [HBM+21] to $\mathcal{A}$ with the bitstring $0^{\lfloor 1/\epsilon\rfloor}$ fixed as input we obtain a polynomial-time classical algorithm $\mathcal{A}^{\prime}$ with the following property: for every $n$ there exists an advice string $\alpha_{n,\epsilon}\in\{0,1\}^{\mathrm{poly(n,1/\epsilon)}}$ [^1] Note that the advice string also depends on $\epsilon$, since $0^{\lfloor 1/\epsilon\rfloor}$ was fixed as input to $\mathcal{A}$. such that for every $x\in I^{n}_{\text{correct}}(\epsilon)$:
@@
Appendix B Discrete logarithm concept class
B.1 $\mathsf{CC}/\mathsf{QQ}$ separation from intractability assumption in [BM84]
In this section we show how to construct a binary concept class that achieves a $\mathsf{CC/QQ}$ separation when one assumes the classical intractability of the discrete logarithm class as stated in [BM84]. More precisely, in [BM84] they state that if a family of classical circuits $\{C_{n}\}$ satisfies
$C_{n}(x,g,p)=\log_{(a,p)}(x)\quad\forall a\in\{1,\dots,p-1\}\text{ and }\forall x\in\mathbb{Z}^{*}_{p}.$ (24)
for at least a $\frac{1}{\mathrm{poly}(n)}$ fraction of all primes $p$, then the size of $C_{n}$ grows faster than any polynomial in $n$.
Remark.
$\log_{(a,p)}x$ denotes the discrete logarithm modulo $p$ of $x$ with respect to the generator $a$. That is, the discrete logarithm $\log_{(a,p)}x$ is the smallest positive integer $\ell$ such that $a^{\ell}\equiv x\mod p.$
See 15
See 2
Proof.
For quantum learnability, we note that a quantum learner can with high probability recover the tuple $(a,p)\in\{0,1\}^{2n}$ from $\mathcal{O}\left(\mathrm{poly}(n)\right)$ examples drawn from $E(c_{(a,p)},\mathcal{D}_{n}^{U})$ (i.e., it just considers those for which $b=1$). Once the quantum learner has figured out what $(a,p)$ is, it can directly use Shor’s algorithm [Sho99] to compute the concept $c_{(a,p)}$ on unseen datapoints $(b,j,x)\in\mathcal{X}_{n}$.
For classical non-learnability, we note that since the examples are efficiently generatable classically, the existence of a classical learner implies the existence of poly($n$)-sized classical circuits $C_{n}$ that satisfy
$\mathsf{Pr}_{(b,j,x)\sim\mathcal{D}_{n}^{U}}\left[C_{n}(a,p,b,j,x)=c_{(a,p)}(b,j,x)\right]\geq 1-\epsilon,\quad\text{for some }\epsilon\in\Omega(1/\mathrm{poly}(n)).$ (25)
In particular, the circuit $C_{n}$ generates examples for $c_{(a,p)}$ based on its input $(a,p)$, runs the classical learning algorithm and afterwards evaluates the hypothesis output by the learner on $(b,j,x)$. What remains to be shown is that this circuit $C_{n}$ can be used to compute $\mathds{1}_{[0,(p-1)/2]}(\log_{(a,p)}(x))$ on at least a $\frac{1}{2}+\frac{1}{\mathrm{poly}(n)}$ fraction of $x\in\mathbb{Z}_{p}^{*}$, since by the random self-reducibility outlined in [BM84] this would violate the hardness of the discrete logarithm stated in [BM84]. To see why this holds, note that the worst $C_{n}$ can do for computing $\mathds{1}_{[0,(p-1)/2]}(\log_{(a,p)}(x))$ while satisfying Eq. (25) is when it is correct on all inputs with $b=1$, and if it is correct on at least one $j^{\prime}$ for a given $x^{\prime}$ then it must be correct on all $j=1,\dots,n$ for that $x^{\prime}$. However, note that even in this case $C_{n}$ can correctly compute $\mathds{1}_{[0,(p-1)/2]}(\log_{(a,p)}(x))$ on at least a $1-2\epsilon$ fraction of all possible $x$. Finally, as $\epsilon\in\Omega(1/\mathrm{poly}(n))$, we conclude that $C_{n}$ can compute $\mathds{1}_{[0,(p-1)/2]}(\log_{(a,p)}(x))$ on at least a $\frac{1}{2}+\frac{1}{\mathrm{poly}(n)}$ fraction of $x\in\mathbb{Z}_{p}^{*}$, thereby violating the hardness of the discrete logarithm stated in [BM84].
B.2 Singleton $\mathsf{CC}/\mathsf{QQ}$ separation from intractability assumption in [BM84]
In this section we show how to construct a non-binary singleton concept class that achieves a $\mathsf{CC/QQ}$ separation when one assumes the classical intractability of the discrete logarithm class as stated in [BM84] (see previous section for the formal statement of intractability stated in [BM84]).
See 16
@@ B.2 Singleton $\mathsf{CC}/\mathsf{QQ}$ separation from intractability assumption in [BM84]
For quantum learnability, we note that the quantum learner does not need data, as it can output the hypothesis that reads inputs $(a,p,x)$ and uses Shor’s algorithm to compute $\log_{(a,p)}(x)$.
For classical non-learnability, we first recap the following claims about “random self-reducibility”.
Claim 1: Fix a prime $p$, due to “random self-reducibilities” we have
  • For any fixed $a$, any polynomial-time algorithm $\mathcal{A}$ that satisfies
    $\mathsf{Pr}_{x\in_{U}\mathbb{Z}_{p}^{*}}\left(\mathcal{A}(a,p,x)=\log_{(a,p)}(x)\right)\geq\frac{1}{2}+\frac{1}{\mathrm{poly}(n)}$
    can be turned into a randomized polynomial-time algorithm $\mathcal{A}^{\prime}$ that satisfies
    $\mathcal{A}^{\prime}(a,p,x)=\log_{(a,p)}(x),\quad\forall x\in\mathbb{Z}^{*}_{p}.$
    proof of claim: $\mathcal{A}^{\prime}$ on input $(a,p,x)$ samples a uniformly random $c\in\{0,\dots,p-1\}$ and computes $b=\mathcal{A}(a,p,xa^{c})$. If $a^{b-c}=x$, output $b-c$. Otherwise, resample $c\in\{0,\dots,p-1\}$ and repeat.
  • Any polynomial-time algorithm $\mathcal{A}$ that satisfies
    $\mathsf{Pr}_{a\in_{U}\mathbb{Z}_{p}^{*}}\Big{(}\forall x\in\mathbb{Z}_{p}^{*}\text{ }:\text{ }\mathcal{A}(a,p,x)=\log_{(a,p)}(x)\Big{)}\geq\frac{1}{2}+\frac{1}{\mathrm{poly}(n)}$
    can be turned into an polynomial-time algorithm $\mathcal{A}^{\prime}$ that satisfies
    $\mathcal{A}^{\prime}(a,p,x)=\log_{(a,p)}(x),\quad\forall a\in\mathbb{Z}^{*}_{p}.$
    proof of claim: $\mathcal{A}^{\prime}$ on input $(a,p,x)$ samples a uniformly random $a^{\prime}\in\{0,\dots,p-1\}$ and computes $b=\mathcal{A}(a^{\prime},p,a)$ and $c=\mathcal{A}(a^{\prime},p,x)$. If $\mathcal{A}$ is correct on $a^{\prime}$, i.e., we have that
    $(a^{\prime})^{b}=a,\text{ }(a^{\prime})^{c}=x\text{ and }\gcd(b,p-1)=1,$
    then output $b^{-1}c$. Otherwise, resample $a^{\prime}\in\{0,\dots,p-1\}$ and repeat.
Claim 2: Fix a prime $p$, a result of Claim 1 is that any polynomial-time algorithm $\mathcal{A}$ that satisfies
$\mathsf{Pr}_{(a,x)\in_{U}\left(\mathbb{Z}_{p}^{*}\right)^{2}}\Big{(}\mathcal{A}(a,p,x)=\log_{(a,p)}(x))\Big{)}\geq\frac{3}{4}+\frac{1}{P(n)}$ (26)
for a polynomial $P$ can be turned into a polynomial-time algorithm $\mathcal{A}^{\prime}$ that satisfies
$\mathcal{A}^{\prime}(a,p,x)=\log_{(a,p)}(x)\quad\forall a\in\{1,\dots,p-1\}\text{ and }\forall x\in\mathbb{Z}^{*}_{p}.$ (27)
proof of claim: Let $P$ be a polynomial, and suppose $\mathcal{A}$ satisfies Eq. (26). We say “$\mathcal{A}$ solves $a$” if
$\mathsf{Pr}_{x\in_{U}\mathbb{Z}_{p}^{*}}\left(\mathcal{A}(a,p,x)=\log_{(a,p)}(x)\right)>\frac{1}{2}+\frac{1}{P(n)-2}.$
That is, the number of tuples $(a,x)$ such that $\mathcal{A}(a,p,x)$ is incorrect is at most $(\frac{1}{2}-\frac{1}{P(n)-2})\cdot 2^{n}$
The question we now ask ourselves is: what is the smallest amount of $a$ that $\mathcal{A}$ can solve while satisfying Eq. (26)? The number of $a$ that $\mathcal{A}$ cannot solve is maximized if $\mathcal{A}$ barely fails to solve it on all the $a$ that it cannot solve. Specifically, for every $a$ that $\mathcal{A}$ cannot solve we have
$\mathsf{Pr}_{x\in_{U}\mathbb{Z}_{p}^{*}}\left(\mathcal{A}(a,p,x)=\log_{(a,p)}(x)\right)=\frac{1}{2}+\frac{1}{P(n)-2}.$
That is, for every $a$ that $\mathcal{A}$ cannot solve the number of tuples $(a,x)$ such that $\mathcal{A}(a,p,x)$ is incorrect is $(\frac{1}{2}-\frac{1}{P(n)}-2)2^{n}$. Let $f$ denote the number of $a$ that $\mathcal{A}$ cannot solve, from Eq. (26) we get
$f\cdot\left(\frac{1}{2}-\frac{1}{P(n)-2}\right)2^{n}<\left(\frac{1}{4}-\frac{1}{P(n)}\right)2^{2n}.$
which shows that $f<\left(\frac{1}{2}-\frac{1}{P(n)}\right)2^{n}$.
Part 1 of Claim 1 implies that if $\mathcal{A}$ solves $a$, then $\mathcal{A}$ can be turned into a polynomial-time randomized algorithm $\mathcal{A}^{\prime}$ that is correct on all $x$ for the given generator $a$. Finally, since the number of $a$ that $\mathcal{A}$ cannot solve is less than $\left(\frac{1}{2}-\frac{1}{P(n)}\right)2^{n}$, part 2 of Claim 1 implies that $\mathcal{A}^{\prime}$ can be turned into a polynomial-time randomized algorithm $\mathcal{A}^{\prime\prime}$ that satisfies
$\mathcal{A}^{\prime\prime}(a,p,x)=\log_{(a,p)}(x)\quad\forall a\in\{1,\dots,p-1\}\text{ and }\forall x\in\mathbb{Z}^{*}_{p}.$ (28)
Using the above two claims, we are able to show that no classical learning algorithm can achieve “efficient learnability”. In particular, the existence of an efficient classical learner implies the existence of a polynomial–time algorithm $\mathcal{A}$ that satisfies
$\mathsf{Pr}_{(a,p,x)}\left(\mathcal{A}(a,p,x)=\log_{(a,p)}(x)\right)\geq 1-\frac{1}{\mathrm{poly}(n)}$ (29)
which we can show – using Claims 1 and 2 – can be turned into an algorithm $\mathcal{A}^{\prime}$ that satisfies
$\mathcal{A}^{\prime}(a,p,x)=\log_{(a,p)}(x)\quad\forall a\in\{1,\dots,p-1\}\text{ and }\forall x\in\mathbb{Z}^{*}_{p}$ (30)
for more than a polynomial fraction of all primes (violating the intractability as outlined in [BM84]).
To see this, we suppose that an algorithm $\mathcal{A}$ satisfies Eq. (29), and show that it will satisfy Eq. (26) for at least a $1-1/\mathrm{poly}(n)$ fraction of all primes, which violates the intractability of the discrete logarithm as outlined in [BM84]. We do so by letting $E$ denote the set of triples $(a,p,x)$ on which $\mathcal{A}$ is incorrect and asking ourselves what the most “prime costly” way for $\mathcal{A}$ to fail (note here we say that $\mathcal{A}$ fails on some prime if it does not satisfy Eq. (26) for this prime). Here by “prime costly” we mean: what is the cheapest way (in terms of number of triples in $E$) for $\mathcal{A}$ to fails on a given prime $p^{\prime}$. In other words, what is the largest amount of primes on which $\mathcal{A}$ can fail.
In short, we note that the most “prime costly” would be to barely fail on every prime $p$ on which it fails. That is, for every prime $p^{\prime}$ on which $\mathcal{A}$ fails, we have that
#$\#\{(a,p^{\prime},x)\in E\}=\left(\frac{3}{4}+\frac{1}{\mathrm{poly(n)}}\right)\cdot 2^{2n}$
Finally, this leads us to the observation that
[…]
where we use that $|E|\leq\left(1-\frac{1}{\mathrm{poly(n)}}\right)\cdot 2^{3n}$.
Appendix C Discrete cube root concept class
C.1 $\mathsf{CC}/\mathsf{QC}$ separation from intractability assumption in [KV94b]
In this section we show how to construct a binary concept class that achieves a $\mathsf{CC/QC}$ separation when one assumes the classical intractability of the discrete cube root as stated in [KV94b]. More precisely, in [KV94b] they state that if a family of classical circuits $\{C_{n}\}$ satisfies
$C_{n}(N,x)=f_{N}^{-1}(x)\quad\forall x\in\mathbb{Z}^{*}_{N}.$ (31)
for at least a $\frac{1}{\mathrm{poly}(n)}$ fraction of moduli $N$, then the size of $C_{n}$ grows faster than any polynomial in $n$.
See 18
See 5
Proof.
For quantum learnability, we note that a quantum learner can with high probability recover the modulus $N$ from $\mathcal{O}\left(\mathrm{poly}(n)\right)$ examples drawn from $E(c_{(N,i)},\mathcal{D}_{n}^{U})$ (i.e., it just considers those for which $b=1$). Once the quantum learner has figured out what $N$ is, it can directly use Shor’s algorithm [Sho99] to compute the $d^{*}$ such that $f_{N}^{-1}(x)\equiv x^{d^{*}}\mod N$.
For classical non-learnability, we note that since the examples are efficiently generatable classically, the existence of a classical learner implies the existence of poly($n$)-sized classical circuits $C_{n}$ that satisfy
$\mathsf{Pr}_{(b,j,x)\sim\mathcal{D}_{n}^{U}}\left[C_{n}(N,i,b,j,x)=c_{(N,i)}(b,j,x)\right]\geq 1-\epsilon,\quad\text{for some }\epsilon\in\Omega(1/\mathrm{poly}(n)).$ (32)
In particular, the circuit $C_{n}$ generates examples for $c_{(N,i)}$ based on its input $(N,i)$, runs the classical learning algorithm and afterwards evaluates the hypothesis output by the learner on $(b,j,x)$. What remains to be shown is that this circuit $C_{n}$ can be used to compute $\mathrm{bin}(f_{N}^{-1}(x),n)$ on at least a $\frac{1}{2}+\frac{1}{\mathrm{poly}(n)}$ fraction of $x\in\mathbb{Z}_{p}^{*}$, since by the random self-reducibility outlined in [ACGS88, GMT82] this would violate the hardness of the discrete logarithm stated in [KV94b]. To see why this holds, note that the worst $C_{n}$ can do for computing $\mathrm{bin}(f_{N}^{-1}(x),n)$ while satisfying Eq. (32) is when it is correct on all inputs with $b=1$, and if it is correct on at least one $j^{\prime}$ for a given $x^{\prime}$ then it must be correct on all $j=1,\dots,n$ for that $x^{\prime}$. However, note that even in this case $C_{n}$ can correctly compute $\mathrm{bin}(f_{N}^{-1}(x),n)$ on at least a $1-2\epsilon$ fraction of all possible $x$. Finally, as $\epsilon\in\Omega(1/\mathrm{poly}(n))$, we conclude that $C_{n}$ can compute $\mathrm{bin}(f_{N}^{-1}(x),n)$ on at least a $\frac{1}{2}+\frac{1}{\mathrm{poly}(n)}$ fraction of $x\in\mathbb{Z}_{N}^{*}$, thereby violating the hardness of the discrete logarithm stated in [KV94b].
C.2 Singleton $\mathsf{CC}/\mathsf{QQ}$ separation from intractability assumption in [KV94b]
In this section we show how to construct a binary singleton concept class that achieves a $\mathsf{CC/QQ}$ separation when one assumes the classical intractability of computing the function $f_{N}^{-1}$ as outlined in [KV94b] (see the previous section for the formal statement of intractability stated in [KV94b]).
See 19
See 6
Proof.
For quantum learnability, we note that the quantum learner does not need data, as it can output the hypothesis that reads inputs $(N,x)$ and uses Shor’s algorithm to compute $f_{N}^{-1}(x)$.
For classical non-learnability, we note that since the examples are efficiently generatable classically, the existence of an efficient classical learner implies the existance of a polynomial-time classical algorithm $\mathcal{A}$ that satisfies
$\mathsf{Pr}_{x\in\mathbb{Z}_{N}^{*}}\left[\mathcal{A}(N,x)=\mathrm{bin}(f^{-1}_{N}(x),n)\right]\geq\frac{1}{2}+\frac{1}{\mathrm{poly}(n)}.$ (33)
for at least an inverse polynomial fraction of moduli $N$. Since computing the least-significant bit of $f_{N}^{-1}$ on a $\frac{1}{2}+\frac{1}{\mathrm{poly}(n)}$ fraction of $x\in\mathbb{Z}_{N}^{*}$ is as hard as to computing $f^{-1}_{N}$ on all $x\in\mathbb{Z}_{N}^{*}$ [ACGS88, GMT82], we conclude that $\mathcal{A}$ can be turned into a polynomial-time algorithm that computes $f^{-1}_{N}$ for an inverse polynomial fraction of moduli, violating the hardness assumption outlined in [KV94b]. ∎
Appendix D Proof of Theorem 7
See 7
Proof.